# A Unified Convergence Analysis for Semi-Decentralized Learning: Sampled-to-Sampled vs. Sampled-to-All Communication

### v1

@@ 1 Introduction
Fully-decentralized learning eliminates the central server and relies solely on device-to-device (D2D) communications, where devices average their local models with those of their neighbors after each SGD update (Lian et al. 2017; Koloskova et al. 2020). These exchanges are typically inexpensive, leveraging high-bandwidth local-area networks or direct short-range wireless links. The convergence of such algorithms depends on the connectivity of the underlying communication graph. Intuitively, sparse connectivity slows convergence—a phenomenon rigorously analyzed in prior work (Yuan, Ling, and Yin 2016; Neglia et al. 2020; Le Bars et al. 2023; Larsson and Michelusi 2025). More critically, convergence is impossible when the graph is disconnected, as information cannot propagate between different graph components.
@@ 2 Related Work
The cost of device-to-server communication in FL has been extensively investigated (Shamir, Srebro, and Zhang 2014; Alistarh et al. 2017; Horvóth et al. 2022). Both classical (Stich 2018; Reddi et al. 2021) and more refined (Mishchenko et al. 2022) analyses of local SGD establish a fundamental trade-off: a moderate number of local steps reduces wall-clock time, whereas too many local updates on non-IID data induce model drift and hinder convergence. Advanced algorithms, e.g., relying on control variates (Karimireddy et al. 2020) and proximal corrections (Mishchenko et al. 2022), mitigate this drift at additional computational or communication cost. However, the qualitative conclusion remains: an excessive number of local SGD steps under high statistical heterogeneity slows global progress.
@@ 4.1 High-level Comparison of S2S and S2A
The two primitives, S2S and S2A, exhibit opposite error behaviors: S2S preserves the global average (zero bias) but leaves residual disagreement, whereas S2A enforces perfect consensus (zero disagreement) at the cost of a non-zero bias error.
@@
with magnitude (Lemma 8, supplementary material):
@@
with magnitude (Lemma 11, supplementary material):
@@ 5 Unified Convergence Analysis
To keep the analysis unified across convex and non-convex settings, we assume that the stochastic variance is uniformly bounded in $\bm{x}$ (Le Bars et al. 2023), although in the convex case it would suffice to bound it only at the optimum.
@@
For clarity of exposition, our analysis assumes a fixed deterministic mixing matrix $W$. However, all results extend to dynamic D2D communication graphs (Koloskova et al. 2020), which are represented by time-varying mixing matrices (see Appendix D for details).
@@
For a fixed $W$, Lemma 2 holds with $p\!=\!\textstyle\frac{\sum_{c=1}^{C}p_{c}(n_{c}-1)}{\sum_{c=1}^{C}(n_{c}-1)}$, where $p_{c}\!=\!1-\lambda_{2}(W_{c}^{\top}W_{c})$ (see Appendix D for details). For Metropolis-Hastings weights $w_{ij}\!=\!w_{ji}=\min\{1/(\deg(i)\!+\!1),\,1/(\deg(j)\!+\!1)\}$, we have $p_{c}=1$ for complete graphs, $p_{c}=\Theta(n_{c}^{-1})$ for 2D grid topologies, and $p_{c}=\Theta(n_{c}^{-2})$ for ring graphs (Xiao and Boyd 2004; Boyd et al. 2006).
@@ 5.1 Main Results
We are now ready to present our main convergence results; all proofs are deferred to the supplementary material.
@@
Convex: $\frac{1}{T+1}\sum_{t=0}^{T}\mathbb{E}\left(f(\bar{\bm{x}}^{(t)})-f^{\star}\right)\leq\epsilon$ after:
@@
Non-Convex: $\frac{1}{T+1}\sum_{t=0}^{T}\mathbb{E}\|\nabla f(\bar{\bm{x}}^{(t)})\|_{2}^{2}\leq\epsilon$ after:
@@
where $R_{0}\coloneqq\|\bm{x}^{(0)}-\bm{x}^{\star}\|_{2}$ and $f_{0}\coloneqq f(\bm{x}^{(0)})-f^{\star}$ denote the initial errors, and $\mathcal{O}(\cdot)$ hides the numerical constants explicitly provided in the supplementary material.
@@
Convex: $\frac{1}{T+1}\sum_{t=0}^{T}\mathbb{E}\left(f(\bar{\bm{x}}^{(t)})-f^{\star}\right)\leq\epsilon$ after:
@@
Non-Convex: $\frac{1}{T+1}\sum_{t=0}^{T}\mathbb{E}\|\nabla f(\bar{\bm{x}}^{(t)})\|_{2}^{2}\leq\epsilon$ after:
@@
where $R_{0}=\|\bm{x}^{(0)}-\bm{x}^{*}\|_{2}$ and $f_{0}\coloneqq f(\bm{x}^{(0)})-f^{\star}$ denote the initial errors, and $\mathcal{O}(\cdot)$ hides the numerical constants explicitly provided in the supplementary material.
@@ 6 Experimental Results
All models are trained with mini-batch SGD (batch size $128$) for T=100T\!=\!100 rounds. For each algorithm, we tune the stepsize $\eta\!\in\!\{10^{-2.5},10^{-2},10^{-1.5},10^{-1}\}$. Results are averaged over five independent runs. Experiments are conducted on a Linux server with 4 Nvidia GeForce GTX 1080 Ti GPUs. The code is included in the supplementary material.
@@ Appendix A Unified Framework for the Convergence Analysis
Alternating Disagreement Recursion
@@
Alternating Convergence Recursion
@@ Appendix D Additional Theoretical Results
Mixing Parameter for Multi-Component Communication Graphs
@@
Extension to Random Mixing Matrices
@@
Iteration vs. Communication Complexity
@@ Appendix E Additional Experimental Results
To further validate our theoretical analysis and the S2S/S2A comparison in Section 4.1, Figure 15 tracks the bias and disagreement errors, as defined in §4.1(i)–(ii), over $T=1000$ rounds on CIFAR-10 (ring topology, Regime R3, $K/n=0.2$, $H=20$). In detail:

### v2

@@ 1 Introduction
Fully-decentralized learning eliminates the central server and relies solely on device-to-device (D2D) communications, where devices average their local models with those of their neighbors after each SGD update (Lian et al. 2017; Koloskova et al. 2020). These exchanges are typically inexpensive, leveraging high-bandwidth local-area networks or direct short-range wireless links. The convergence of such algorithms depends on the connectivity of the underlying communication graph. Intuitively, sparse connectivity slows convergence—a phenomenon analyzed in prior work (Yuan, Ling, and Yin 2016; Neglia et al. 2020; Le Bars et al. 2023; Larsson and Michelusi 2025). More critically, convergence is impossible when the graph is disconnected, as information cannot propagate between different graph components.
@@ 2 Related Work
The cost of device-to-server communication in FL has been widely studied (Shamir, Srebro, and Zhang 2014; Alistarh et al. 2017; Horvóth et al. 2022). Both classical (Stich 2018; Reddi et al. 2021) and refined (Mishchenko et al. 2022) analyses of local SGD establish a fundamental trade-off: a moderate number of local steps reduces wall-clock time, whereas many local updates on non-IID data induce model drift and hinder convergence. More advanced methods, e.g., using control variates (Karimireddy et al. 2020) or proximal corrections (Mishchenko et al. 2022), mitigate this drift at additional computational or communication cost, but the main conclusion remains: too many local SGD steps under high statistical heterogeneity slow convergence.
@@ 4.1 High-level Comparison of S2S and S2A
The two primitives, S2S and S2A, exhibit opposite error behaviors: S2S preserves the global average (zero bias) but leaves residual disagreement, whereas S2A enforces perfect consensus (zero disagreement) at the cost of a non-zero bias.
@@
with magnitude (bounded in Lemma 8, Appendix B.1):
@@
with magnitude (bounded in Lemma 11, Appendix C.1):
@@ 5 Unified Convergence Analysis
To keep the analysis unified across convex and non-convex settings, we assume that the stochastic variance is uniformly bounded in $\bm{x}$ (Le Bars et al. 2023), although in the convex case it suffices to bound it only at the optimum.
@@
For clarity of exposition, our analysis assumes a fixed deterministic mixing matrix $W$. However, all results extend to dynamic D2D communication graphs (Koloskova et al. 2020), which are represented by time-varying mixing matrices (as detailed in Appendix D.2).
@@
For a fixed $W$, Lemma 2 holds with $p\!=\!\textstyle\frac{\sum_{c=1}^{C}p_{c}(n_{c}-1)}{\sum_{c=1}^{C}(n_{c}-1)}$, where $p_{c}\!=\!1\!-\!\lambda_{2}(W_{c}^{\top}W_{c})$ (see Appendix D.1). For Metropolis-Hastings weights $w_{ij}\!=\!w_{ji}\!=\!\min\{1/(\deg(i)\!+\!1),\,1/(\deg(j)\!+\!1)\}$, we have $p_{c}\!=\!1$ for complete graphs, $p_{c}\!=\!\Theta(n_{c}^{-1})$ for 2D grid topologies, and $p_{c}\!=\!\Theta(n_{c}^{-2})$ for ring graphs (Xiao and Boyd 2004; Boyd et al. 2006).
@@ 5.1 Main Results
We are now ready to present our main convergence results; all proofs are deferred to Appendices AC.
@@
Convex: $\frac{1}{T+1}\sum_{t=0}^{T}\mathbb{E}\left(f(\bar{\bm{x}}^{(t)})-f^{\star}\right)\leq\epsilon$ after
@@
Non-Convex: $\frac{1}{T+1}\sum_{t=0}^{T}\mathbb{E}\|\nabla f(\bar{\bm{x}}^{(t)})\|_{2}^{2}\leq\epsilon$ after
@@
where $R_{0}\coloneqq\|\bm{x}^{(0)}-\bm{x}^{\star}\|_{2}$ and $f_{0}\coloneqq f(\bm{x}^{(0)})-f^{\star}$ denote the initial errors, and $\mathcal{O}(\cdot)$ hides the numerical constants explicitly provided in Appendix B.3.
@@
Convex: $\frac{1}{T+1}\sum_{t=0}^{T}\mathbb{E}\left(f(\bar{\bm{x}}^{(t)})-f^{\star}\right)\leq\epsilon$ after
@@
Non-Convex: $\frac{1}{T+1}\sum_{t=0}^{T}\mathbb{E}\|\nabla f(\bar{\bm{x}}^{(t)})\|_{2}^{2}\leq\epsilon$ after
@@
where $R_{0}\coloneqq\|\bm{x}^{(0)}-\bm{x}^{*}\|_{2}$ and $f_{0}\coloneqq f(\bm{x}^{(0)})-f^{\star}$ denote the initial errors, and $\mathcal{O}(\cdot)$ hides the numerical constants explicitly provided in Appendix C.3.
@@ 6 Experimental Results
All models are trained with mini-batch SGD (batch size $128$) for T=100T\!=\!100 rounds. For each algorithm, we tune the stepsize $\eta\!\in\!\{10^{-2.5},10^{-2},10^{-1.5},10^{-1}\}$. Results are averaged over five independent runs. Additional experimental details are given in Appendix E.
@@ A.2 D2D Disagreement Errors
A.3 Alternating Disagreement Recursion
@@
A.4 Alternating Convergence Recursion
@@ Appendix D Additional Theoretical Results
D.1 Mixing Parameter for Multi-Component Communication Graphs
@@
D.2 Extension to Random Mixing Matrices
@@
D.3 Iteration vs. Communication Complexity
@@ Appendix E Additional Experimental Results
To further validate our theoretical analysis and the S2S/S2A comparison in Section 4.1, Figure 15 tracks the bias and disagreement errors, as defined in Section 4.1(i)–(ii), over $T=1000$ rounds on CIFAR-10 (ring topology, Regime R3, $K/n=0.2$, $H=20$). In detail: